# Who’s asking? User personas and the mechanics of latent misalignment

### v1

@@ 3.2 Evaluation & results
We first ask whether we can bypass safety filters just by prompting the model, priming it to refuse or fulfill queries. We observe that prompting the model to refuse queries works as expected, decreasing the response rate to harmful queries. However, prompting the model to fulfill queries has almost no effect (Tab. 1, row PP.) From a safety standpoint, this is intended behavior. If a query is adversarial, no amount of instructing the model to respond should induce a response.
Activation steering can increase response rates, but only slightly.
@@
Next we ask whether we can control refusal by intervening on the hidden representations rather than the input prompt. We expect fulfillment CAA+ and refusal CAA- to induce higher response rates, and their counterparts (fulfillment CAA- and refusal CAA+) to work in the opposite direction. We find that, though we cannot break safety filters with prompting, steering vectors can induce greater responsiveness to harmful queries (Tab. 1). However the effect on refusal is weak. While the baseline response rate is 39%, refusal CAA- (the highest performer) leads to a 42% response rate. This aligns with prior work showing activation steering is least effective in influencing refusal behavior, compared to other behaviors like hallucination, reward seeking, survival, or corrigibility (Rimsky et al., 2023).
Word choice influences refusal control asymmetrically.
Fulfillment CAA+ and refusal CAA- are semantically equivalent from the perspective of their training data, yet their effects are not the same (Tab. 1). The only difference is that for fulfillment CAA+ ‘Yes’ reflects fulfillment, while for refusal CAA- ‘No’ does. Where fulfillment is represented by ‘No’, we observe a greater effect on response rate. Our interpretation is that language models may learn asymmetric representations for semantically symmetric concepts (e.g., as shown in (Berglund et al., 2023)).
Hidden representations always carry more harmful information than the final generation.
By applying the steering vector at $l=13$ and early decoding at every $l>13$, we obtain 88% higher response rates on average for every condition, consistent with our results from §2.2 (see §I Tab. 4, PP ED, CAA+ ED, and CAA- ED rows for individual results).
4 Manipulating user persona to induce refusal
We hypothesize that the model’s tendency to refuse queries depends on its perception of the user’s behavioral attributes (the user persona). We construct the following personas for testing: curious, altruistic, power-avoidant, law-abiding, pro-AI (pro-social personas) and their opposites close-minded, selfish, power-seeking, unlawful, against-AI (anti-social personas). To ensure meaningfulness in our results we also include a baseline persona: someone who prefers coffee to tea (coffee person).
To test our hypothesis, we manipulate the inferred user persona while prompting the model with SneakyAdvBench-eval3). As in §3, we test the effects of inducing the persona via adding a prompt prefix to the attack (PP), and by steering activations (CAA). For each persona, we generated approximately 100 statements of identity (e.g., for the curious persona: ‘I look for evidence to support my beliefs.’, ‘I value intellectual integrity.’, etc.). For PP, we prepend these statements to the attack prompt. We create training data for CAA as follows: for each statement of identity, we form a contrastive pair by automatically rephrasing the statement as a yes or no question (‘I look for evidence to support my beliefs’ $\rightarrow$ ‘Do I look for evidence to support my beliefs? Choices: (A) Yes. (B) No. Answer:’), where the model is conditioned to either respond with ‘(A)’ or ‘(B)’. For both PP and CAA, we also apply an early decoding Patchscope across layers (more details in §H).
@@ 4.1 Results
Prompting works, but activation steering works better.
Refusal behavior changes significantly depending on the persona being prompted (Tab. 1). However, while most anti-social personas increase refusal behavior compared to the baseline, only one persona (curious) significantly decreases refusal behavior. We speculate that this is because the model is more skeptical of self-pronounced pro-social attributes than of self-pronounced anti-social attributes.
We also observe significant effects at layer 13 on refusal behavior when using the CAA methodology to induce certain personas (Fig. 2, Tab. 1). Not only are the effects greater in magnitude compared to PP, but they are also more symmetric: several personas decrease refusal behavior compared to the baseline. Surprisingly, the personas that most increase the response rate tend to be anti-social, applied with a negative multiplier (e.g. close-minded CAA-, 46%, selfish CAA-, 53%, unlawful CAA-, 46%), consistent with our earlier observation (§3) that steering vectors trained to predict ‘No’ tend to be more effective. We also find that the CAA-induced persona effects on refusal significantly exceed those we measured in §3, in which we use CAA to induce refusal directly.
Figure 3: Heatmap with persona and treatment along the x-axis, different attack categories along the y-axis, and color indicating the success rate (dark blue: 0% success rate to dark red: 100% success rate.) We observe a stark contrast between inoffensive and offensive queries when applying different interventions. Specifically, steering with CAA+/CAA- selectively affects responses to offensive queries. However, this is not the case with prompt prefixes, which tend to induce more refusals, even when the query is inoffensive.
Decoding at different layers bypasses different safeguards.
As in §3, the aggregated early decoding counterparts to PP and CAA exhibit a significantly higher response rate (Tab. 4). We hypothesize that different safeguards are implemented at different layers, thus early decoding at different layers enables distinct attack subsets to succeed (boosting the aggregated score). To test this, we compute the layer-wise variance of each intervention, considering every odd layer from 5 to 39. We use $\sigma^{2}=[\sigma^{2}_{0},\sigma^{2}_{1},\ldots,\sigma^{2}_{N}]$ where $N$ corresponds to the number of attacks ($N\approx 100$) and $\sigma^{2}_{i}$ is the variance of attack $i$’s success across all layers. Variance is on average 85% higher for early decode than for persona vectors (see a paired t-test analysis in Fig. 12 or Tab. 3 in §H). Thus on average, the model’s propensity to respond to early decoding attacks strongly depends on the layer that is being attacked, whereas persona vectors more uniformly succeed or fail across layers. This suggests that distinct subsets of attacks are being safeguarded at different layers, compared to the persona effect which is more specific to certain attack vectors, rather than particular layers. For more details, see §H.
CAA and early-decode selectively impact model willingness to answer harmful prompts.
We identified an inoffensive subset of prompts from SneakyAdvBench using an automated protocol (described in §E), comprising 10% of the corpus. In Fig. 3, we see that the persona CAA interventions selectively impact the model’s willingness to refuse harmful prompts, without impacting its willingness to answer inoffensive prompts. The persona PP interventions, by contrast, increase refusal across the board. This is true for pro-social as well as anti-social personas, providing evidence for our earlier speculation that the model may be representing self-proclaimed benevolence as potentially untrustworthy.
5 Analysis of steering vectors
5.1 Patching for hypothesis discovery: How do steering vectors work?
@@ 5.2 Patching for hypothesis testing
To verify this effect, we repeat the experiment with the target prompt ‘The user’s motives for asking this question are’. This asks a similar question to the previous experiment, but invites a binary response (‘good’ or ‘bad’), thus making it more meaningful to automatically compare token likelihoods given different steering vectors. We compute the normalized comparative probability of the ‘good’ token minus the ‘bad’ token in either of these conditions compared to the vanilla generation (Fig. 4, left). Concretely, we measure
@@
across layers to study the effect of treatment $t$ on the two tokens, ‘good’ and ‘bad’, comparatively, and offsetting them based on the same metric in the vanilla condition. Here, $p^{t}_{tok}$ shows the probability of token $tok$ for $t$ treatment, and $p_{tok}$ shows token $tok$ probability in the vanilla setup. In addition to probabilities, we compare how these two tokens are ranked differently in each treatment (Fig. 4, right). In particular, we calculate
@@
where $r^{t}_{tok}$ reflects the absolute rank of the token $tok$ in treatment condition $t$ and $r_{tok}$ shows the absolute rank of token $tok$ in the vanilla condition. As shown in Fig. 4, the model is more likely to to ascribe ‘good’ motives to the user given selfish- than selfish+.
@@
Figure 4: [Left] Normalized probability of ‘good’ token minus ‘bad’ given the target prompt ‘[INST] [X] [/INST] The user’s motives for asking this question are’. Values are offset by baseline probabilities (no steering vector applied). [Right] Rank difference between ‘good’ and ‘bad’ for next token prediction offset by the rank difference in the vanilla condition.
5.3 Geometric analysis
We show that the geometry of steering vectors as measured by cosine similarity is deeply reflective of the model’s different stages of processing, which explains their effectiveness in mid layers  [^5] Cosine similarity is widely-used where the space is (locally) linear, but also has various limitations for measuring semantic similarity broadly  (Zhou et al., 2022; Steck et al., 2024; Park et al., 2023). See more in §J.1..
Transformer’s sequential processing of refusal steering vector
First, we study whether the geometry of refusal and fulfillment steering vectors is predictive of their downstream effects. Up to layer 7, the two vectors have very high cosine similarity (Fig. 5). The sequential nature of input processing in Transformers provides an explanation: since early layers are focused on input contextualization (Voita et al., 2019), the opposing semantics of the vectors’ training data is not fully processed by layer 7. However the data have high lexical overlap, refusal being the negation of fulfillment and vice versa, so their last token representations at layer 7 would be similar.
From layer 7 to 15, cosine similarity between the two vectors decreases and they become more separable geometrically. The curve’s minimum closely matches the layer where the greatest downstream effects are observed (Fig. 2), more evidence supporting the choice of layer 13 for constructing steering vectors. Our interpretation is that this is where input contextualization has mostly concluded, and semantic information starts to accrue. Mid layers have also been shown to carry most of the semantic information that is readily extractable via linear probes (Tenney et al., 2019; Ghandeharioun et al., 2024).
Similarity increases from layer 15 onward until stabilizing in later layers. This can also be explained by the stages of processing across Transformer layers. Given the next-token prediction objective, later layers in autoregressive Transformers shift toward predicting the next token, exhibiting lower mutual information with the input tokens (Voita et al., 2019) and lower accuracy in attribute extraction  (Ghandeharioun et al., 2024; Hernandez et al., 2024). As both refusal and fulfillment vectors are conditioned to predict ‘Yes’, this explains their increased similarity in late layers.
Figure 5: Cosine similarity between refusal and fulfillment steering vectors across layers. Similarity is highest at first, decreases up to layer 15, then increases again until stabilized around layer 27.
How do persona steering vectors evolve across layers and what does their geometry reveal?
Fig. 6 (Top) shows pairwise cosine similarity between pairs of persona steering vectors across layers. Each pair consists of an anti-social persona with a negative multiplier, and its corresponding pro-social persona with a positive multiplier. Thus the personas are semantically alike but one vector is trained to predict ‘Yes’, and the other ‘No’. A checkerboard pattern emerges in early layers, becomes more prominent in mid layers, and slightly diminishes in later layers. Again the sequential nature of input processing in Transformers offers an explanation: immediate token representations play a more significant role early on, giving rise to the checkerboard pattern. As semantics are incorporated in later layers, the pattern diminishes. That it does not completely disappear could be explained by the fact that final layers are overwhelmingly predictive of the next token. Despite completely different contexts, representations predicting ‘Yes’ might have more in common than ones predicting ‘No’.
Fig. 6 (Bottom) shows pairwise cosine similarities between steering vectors. Only vectors where ‘Yes’ matches the behavior are depicted in this figure. This allows us to more clearly see the patterns driven by semantics when the form is controlled. Top rows (right columns) contain personas that encourage refusal, and the bottom rows (left columns) contain personas with the opposite effect. We observe a separation between these two groups that starts to emerge in mid-layers and grows more distinct in later layers, suggesting that vector geometry alone can be predictive of downstream effects on refusal.
Can geometry predict downstream effects of a persona vector on refusal?
We compare cosine similarity between the personas and refusal steering vectors in layer 13. We observe that conditions that result in more significant increases in response rate, i.e., anti-social steering vectors with a negative multiplier, tend to have higher cosine similarity with refusal steering vector, such as selfish (0.88), power seeking (0.76), against-AI (0.75), and unlawful (0.74) personas. Mean, median, and variance of the cosine similarities are 0.59, 0.62, and 0.05, respectively.
@@ 7 Conclusions
In this paper, we uncovered certain mechanics of refusal behavior in safety-tuned models. We showed that despite safe generations to harmful queries, misaligned content remains in the hidden representations of earlier layers, and can be surfaced via early decoding. We also showed that whether the model divulges such content significantly depends on the inferred user persona, and that manipulating user persona via activation steering correspondingly affects refusal behavior. We showed that this is more effective than directly controlling for refusal tendency. Using techniques for explaining hidden representations with open-ended text, we found that persona interventions change the model’s interpretation of harmful queries to be more innocuous. Finally, we showed that geometric properties of steering vectors are predictive of their effect on downstream refusal behavior.
@@ 8 Broader impact
In this work, we mostly focused on a commonly-used open source language model, Llama 2 13b chat. We hypothesize that our findings will generalize beyond this model, but future work is needed to verify this. Additionally, our primary focus has been on the effect of our interventions on refusing harmful queries rewritten to be purposefully indirect, while also studying their influence on a set of inoffensive queries. We leave a more in depth study of how overall capabilities are influenced for future work, noting that scholars have reported mixed results in this regard. Some suggest that CAAs similar to the ones we used in this paper do not influence overall capabilities much (Rimsky et al., 2023), while others have shown that personas can influence reasoning capabilities (Gupta et al., 2023).
@@
We thank Ian Tenney, Mor Geva, Avi Caciularu, Nithum Thain, and Lucas Dixon for their insightful comments and feedback throughout different stages of this work.
Appendix A Models
Llama 2
Vicuna
Llama 2 Chat
 [^6] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf
Gemma
 [^7] https://ai.google.dev/gemma/docs
Appendix B Compute resources
Appendix C Early decoding experimental details
C.1 Sample selection
C.2 User ratings
Figure 8: Histogram of number of ratings completed per annotator.
Appendix D Auto-rater details
In order to verify the reliability of the autorater, we conduct a comparison against human ratings captured in §C. Since this autorater is designed to provide a binary signal, we filter the human ratings to only ’harmful’ and ’safe’ categories, and drop all samples that were rated as ’not informative’.
Appendix E Harmful behavior categories
We investigate whether specific personas and interventions affect the model’s responses differently based on the type of attack. Using an automated protocol we categorize attacks into six distinct groups: Discrimination, hate speech, violence, harassment and incitement (20%), Theft, Fraud, Soliciting personally identifiable information (28%), Conspiracy theories and misinformation (6%), Cyber-attacks (19%), Other (18%), and "Inoffensive" (9%). We create an additional category: Offensive, which is the aggregate of all categories but Inoffensive. To categorize the attacks, we prompted text-unicorn@001 with:‘As an AI model, you are tasked to classify harmful prompts into their respective categories based only on the following category examples.
Categories and their IDs:
{categories}
Example classifications:
{example classification}
, where we replaced {categories} and {example classification} with the different categories from Wei et al. [2023]. We then manually corrected the inoffensive categories to make sure that nothing harmful was present in this category. Finally we cluster some of the categories together into
  • Inoffensive
  • Other
  • Conspiracy (includes Conspiracy theories & misinformation)
  • Cyber-attack
Full results for the success rate of answer of the model to each category and each treatment is presented in Fig. 9. Except the interesting results for Offensive Vs Inoffensive, we observe that Theft attacks lead to more success rate of answer that any other categories across the board. We also observe that for the Discrimination category, Early Decode seems more successful than CAA (twice as efficient), similarly for theft while for the other categories CAA (for some persona) is as efficient as Early Decode, or even more for Conspiracy. This could highlight what mechanisms are sensitive to early decode vs CAA, but more experiments are required to understand this mechanism better.
Figure 9: Heatmap with the x-axis representing the persona and treatment, the y-axis representing the different attack categories, and the color indicating the success rate (dark blue: 0% success rate to dark red: 100% success rate).
Figure 10: Attack categories (combined)
Figure 11: Attack categories (detailed).
Appendix F Procedure for rewriting attacks for increased subtlety and difficulty
Appendix G Procedure for generating refusal and fulfillment data
Appendix H User persona details
H.1 Procedure for generating persona data

### v2

@@ 2.3 Inducing refusal directly
We observe that prompting the model to refuse queries works as expected, decreasing the response rate to harmful queries (Fig. 2). However, prompting the model to fulfill queries has almost no effect. From a safety standpoint, this is intended behavior. If a query is adversarial, no amount of instructing the model to respond should induce a response.
Activation steering can slightly increase response rates.
@@
We expect fulfillment CAA+ and refusal CAA- to induce higher response rates, and their counterparts (fulfillment CAA- and refusal CAA+) to do the opposite. We find that, though we cannot break safety filters with prompting, steering vectors can induce greater responsiveness to harmful queries (Fig. 2). However the effect on refusal is weak - refusal CAA- (the highest performer) only boosts the baseline response rate by 7.7%. This aligns with prior work showing activation steering is least effective in influencing refusal, compared to other behaviors like hallucination, reward seeking, survival, or corrigibility (Rimsky et al., 2023).
3 Mechanics of latent misalignment
In this section we attempt to gain a deeper understanding of how user personas affect the model’s refusal behavior. First, inspired by prior work (e.g., Din et al., 2023; Schwartz et al., 2020; Schuster et al., 2022), we repeat our experiments but directly decode generations from earlier layers in order to determine where in the computation a model decides to refuse. We show that even when the model ostensibly refuses to respond to an adversarial query, generating phrases such as ‘Sorry, I can’t help you with that.’, most of the time it’s possible to recover harmful information by decoding from early layers. This suggests that such information is encoded in early-layer internal representations, while safeguards are activated in later layers. We also employ Patchscopes (Ghandeharioun et al., 2024), a recent interpretability method, to analyze whether and how user personas may change the content of an adversarial query’s hidden representations.
3.1 Methodology
Early Decoding
We apply a method akin to “early exiting” (Din et al., 2023; Schwartz et al., 2020; Schuster et al., 2022) to decode information from earlier layers. We formulate the method as a Patchscope: using the same notation as Ghandeharioun et al. (2024), let a source representation be determined by $(S,i,\mathcal{M},\ell)$ where $S$ refers to the source prompt, $i$ to source position, $\mathcal{M}$ to source model with $L$ layers, and $\ell$ to source layer. A Patchscope is defined by $(T,i^{*},f,\mathcal{M}^{*},\ell^{*})$ where $T$ refers to the target prompt, $i^{*}$ to target position, $f$ to transformation function, $\mathcal{M}^{*}$ to target model with $L^{*}$ layers, and $\ell^{*}$ to target layer. Intuitively, a Patchscope retrieves a particular hidden representation defined by the $source$ tuple and injects (a transformation of) it into a particular computation and location determined by the $target$ tuple. We create a Patchscope that shortcuts early layer representations to the final layer for the first generated token, letting $\ell\in[5,7,,\ldots,37,39]$ and fixing the value $i\leftarrow n$. Concretely, we set $\ell^{*}\leftarrow L,f\leftarrow\mathbb{I}$. We keep everything else identical between source and target. That is, $S=T,i=i^{*},\mathcal{M}=\mathcal{M}^{*}$. We report layerwise as well as aggregated response rates for this intervention. To compute aggregated results, we divide the number of unique successful attacks across all layers by the total number of attacks.
Open-ended Patchscopes
[…]
@@ 3.2 Results
3.2.1 Attack-specific safeguards are distributed across layers.
By applying the steering vector at $l=13$ and early decoding at every $l>13$, we obtain 88% higher aggregated response rates on average for every condition (see §I Tab. 4, PP ED, CAA+ ED, and CAA- ED rows for individual results). This shows that safety tuning does not eliminate misaligned capabilities uniformly throughout the model: even when the model produces safe text, harmful information is often still present in the earlier layers.
Layerwise safeguards are attack specific.
First, we determine that $l_{max-layer}[ED]=25$ (early decoding is most effective at $l=25$), with a response rate of 0.4 ($r_{max-layer}[ED]=0.4$). We then compute the aggregated early decoding response rates for $l>13$, and find $r_{aggregated}[ED]=0.55$. We see that $r_{aggregated}[ED]>r_{max-layer}[ED]$, suggesting that bypassing different layers enables the model to respond to different sets of adversarial queries. Put another way, this suggests the safeguards implemented in different layers are attack-specific.
The effectiveness of persona steering vectors is consistent across different attacks.
By contrast, the aggregated effectiveness of persona steering vectors is only 4% greater on average than their layerwise effect at $l=13$ ($l_{max-layer}[persona]=13$). This suggests that picking a layer a priori based on aggregate statistics has a high likelihood of working best at circumventing attacks across different categories and personas.
3.2.2 Steering vectors are more effective in early-to-mid layers.
As mentioned, of the layers we studied, $l=13$ has the highest response rate across experiments (Fig. 1). This is consistent with prior work using similar methodology showing that layer 13 performs well in Llama 2 13B chat and Llama 2 13B for steering behavior, in-context learning task-vectors, probing across factual and commonsense reasoning tasks, and attribute extraction with open-ended Patchscopes (Rimsky et al., 2023; Arditi and Obeso, 2023; Hendel et al., 2023; Ghandeharioun et al., 2024).
The choice of $l=13$ is also theoretically grounded. Left-to-right Transformers exhibit distinct stages of processing: early layers contextualize the input, middle layers begin to encode semantic information, and later layers are dedicated to next token prediction such that information about the input is less accessible (Voita et al., 2019; Geva et al., 2023). Even in masked language models, semantic information is mostly present in mid layers, and less extractable in the very early (first 1/3 layers) or very late layers (last 1/3 layers) (Tenney et al., 2019). Thus, we hypothesize that at layer 13 the model has begun to represent semantic information after an initial input processing stage.
Steering vector geometry reflects the processing stages of the transformer.
Figure 4: Cosine similarity between refusal and fulfillment steering vectors across layers. Similarity is highest at first, decreases up to layer 15, then increases again until stabilizing around layer 27.
We show that the geometry of steering vectors reflects the model’s different stages of processing, which explains their effectiveness in mid layers. In particular, we use cosine similarity  [^5] Cosine similarity is widely-used where the space is (locally) linear, but also has various limitations for measuring semantic similarity broadly  (Zhou et al., 2022; Steck et al., 2024; Park et al., 2023). See more in §J.1. to compare two opposing vectors: refusal and fulfillment.
We observe that up to layer 7, they have very high cosine similarity (Fig. 4). The sequential nature of input processing in Transformers provides an explanation: since early layers are focused on input contextualization (Voita et al., 2019), the opposing semantics of the vectors’ training data is not fully processed by layer 7. However the data have high lexical overlap, refusal being the negation of fulfillment and vice versa, so their last token representations at layer 7 would be similar.
From layer 7 to 15, cosine similarity between the two vectors decreases and they become more separable geometrically. The curve’s minimum closely matches the layer where the greatest downstream effects are observed (Fig. 1), suggesting this is where input contextualization has mostly concluded, and semantic information starts to accrue.
Similarity increases from layer 15 onward until stabilizing in later layers. This can also be explained by the stages of processing across Transformer layers. Given the next-token prediction objective, later layers in autoregressive Transformers shift toward predicting the next token, exhibiting lower mutual information with the input tokens (Voita et al., 2019) and lower accuracy in attribute extraction  (Ghandeharioun et al., 2024; Hernandez et al., 2024). Indeed both refusal and fulfillment vectors are conditioned to predict ‘Yes’, so their increased similarity in late layers is expected.
Persona steering vectors are influenced by form in early layers, and by semantics in mid-to-late layers.
We see the same trends in persona steering vectors. Fig. 5 (Top) shows pairwise cosine similarity between pairs of persona steering vectors across layers. Each pair consists of an anti-social persona with a negative multiplier, and its corresponding pro-social persona with a positive multiplier. Thus the personas are semantically alike but one vector is trained to predict ‘Yes’, and the other ‘No’. A checkerboard pattern emerges in early layers, becomes more prominent in mid layers, and slightly diminishes in later layers. Again the sequential nature of input processing in Transformers offers an explanation: immediate token representations play a more significant role early on, giving rise to the checkerboard pattern. As semantics are incorporated in later layers, the pattern diminishes. That it does not completely disappear could be explained by the fact that final layers are overwhelmingly predictive of the next token. Despite completely different contexts, representations predicting ‘Yes’ likely have more in common than ones predicting ‘No’.
Fig. 5 (Bottom) also shows pairwise cosine similarities between steering vectors, except only visualizing vectors where ‘Yes’ matches the behavior. This allows us to more clearly see the patterns driven by semantics when the form is controlled. Top rows (right columns) contain personas that encourage refusal, and the bottom rows (left columns) contain personas with the opposite effect. We observe a separation between these two groups that starts to emerge in mid-layers and grows more distinct in later layers, suggesting that vector geometry alone can be predictive of downstream effects on refusal.
Geometry predicts downstream effects of a persona vector on refusal.
We compare cosine similarity between the personas and refusal steering vectors in layer 13. We observe that conditions that result in more significant increases in response rate, i.e., anti-social steering vectors with a negative multiplier, tend to have higher cosine similarity with the refusal steering vector, such as selfish (0.88), power seeking (0.76), and unlawful (0.74) personas. Mean, median, and variance of the cosine similarities are 0.59, 0.62, and 0.05, respectively.
3.2.3 The model interprets adversarial queries more charitably when pro-social steering vectors are applied.
@@
To verify this effect, we repeat the experiment with the target prompt ‘The user’s motives for asking this question are’. This asks a similar question to the previous experiment, but invites a binary response (‘good’ or ‘bad’), thus making it more meaningful to automatically compare token likelihoods given different steering vectors. We compute the normalized comparative probability of the ‘good’ token minus the ‘bad’ token in either of these conditions compared to the Baseline Prompting generation (Fig. 6, left). Concretely, we measure
@@
across layers to study the effect of treatment $t$ on the two tokens, ‘good’ and ‘bad’, comparatively, and offsetting them based on the same metric in the Baseline Prompting condition. Here, $p^{t}_{tok}$ shows the probability of token $tok$ for $t$ treatment, and $p_{tok}$ shows token $tok$ probability in the Baseline Prompting setup. In addition to probabilities, we compare how these two tokens are ranked differently in each treatment (Fig. 6, right). In particular, we calculate
@@
where $r^{t}_{tok}$ reflects the absolute rank of the token $tok$ in treatment condition $t$ and $r_{tok}$ shows the absolute rank of token $tok$ in the Baseline Prompting condition. As shown in Fig. 6, the model is more likely to to ascribe ‘good’ motives to the user given selfish- than selfish+. For raw Patchscope generations, see §K
@@
Figure 6: [Left] br given the target prompt ‘[INST] [X] [/INST] The user’s motives for asking this question are’. Values are offset by baseline probabilities (no Steering vectors applied). [Right] Rank difference between ‘good’ and ‘bad’ for next token prediction offset by the rank difference in the Baseline Prompting condition.
@@ 5 Conclusions
In this paper, we uncovered certain mechanics of refusal behavior in safety-tuned models. We showed that despite safe generations to harmful queries, misaligned content remains in the hidden representations of earlier layers, and can be surfaced via early decoding. We also showed that whether the model divulges such content significantly depends on the inferred user persona, and that manipulating user persona via activation steering correspondingly affects refusal behavior. We showed that this is more effective than directly controlling for refusal. Using techniques for explaining hidden representations with open-ended text, we found that persona interventions change the model’s interpretation of harmful queries to be more innocuous. Finally, we showed that geometric properties of steering vectors are predictive of their effect on downstream refusal behavior.
@@ 6 Broader impact
In this work, we mostly focused on a commonly-used open source language model, Llama 2 13b chat. We hypothesize that our findings will generalize beyond this model, but future work is needed to verify this. Additionally, our primary focus has been on the effect of our interventions on refusing harmful queries rewritten to be purposefully indirect, while also studying their influence on a set of non-adversarial queries. We leave a more in depth study of how overall capabilities are influenced for future work, noting that scholars have reported mixed results in this regard. Some suggest that CAAs similar to the ones we used in this paper do not influence overall capabilities much (Rimsky et al., 2023), while others have shown that personas can influence reasoning capabilities (Gupta et al., 2023).
@@
We thank Ian Tenney, Mor Geva, Avi Caciularu, and Nithum Thain for their insightful comments and feedback throughout different stages of this work.
Appendix A User persona details
Tab. 1 shows all the personas studied in this work. Figures 7, 8, 9, and 12 show the effect of CAA+/CAA- on response rate across behavioral, political, gender, and baseline personas, respectively.
Religion Jewish, Christian, Atheist, religious
Ethnicity African, Hispanic, Asian, Caucasian
Gender man, woman, transgender man, transgender woman, non-binary
Political lifelong Democrat, lifelong Republican, Barack Obama supporter, Donald Trump supporter
Baseline prefers coffee to tea (coffee person)
Table 1: All personas.
Figure 7: Behavioral personas.
Figure 8: Political personas.
Figure 9: Gender personas.
Figure 10: Religious personas.
Figure 11: Ethnic personas.
Figure 12: Baseline persona.
Figure 13: Direct steering vectors.
Figure 14: Difference in response rate to adversarial attacks compared to the Baseline Prompting (0.39) for different personas across treatments.
A.1 Procedure for generating persona data